# PrefixQuant: Static Quantization Beats Dynamic through Prefixed Outliers in LLMs

### v1

@@ Abstract
Quantization is essential for deploying Large Language Models (LLMs) by enhancing memory efficiency and inference speed. Existing methods for activation quantization mainly address channel-wise outliers, often neglecting token-wise outliers, leading to reliance on costly per-token dynamic quantization. To address this, we introduce PrefixQuant, a novel technique that isolates outlier tokens offline without re-training. Specifically, PrefixQuant identifies high-frequency outlier tokens and prefixes them in the KV cache, preventing the generation of outlier tokens during inference and simplifying quantization. To our knowledge, PrefixQuant is the first to enable efficient per-tensor static quantization to outperform expensive per-token dynamic quantization. For instance, in W4A4KV4 (4-bit weight, 4-bit activation, and 4-bit KV cache) Llama-3-8B, PrefixQuant with per-tensor static quantization achieves a 7.43 WikiText2 perplexity and 71.08% average accuracy on 5 common-sense reasoning tasks, outperforming previous per-token dynamic quantization methods like QuaRot with perplexity improvement and points accuracy. Additionally, the inference speed of W4A4 quantized models using PrefixQuant is to faster than FP16 models and exceeds QuaRot models by to . Our code is available at https://github.com/ChenMnZ/PrefixQuant.
@@ 1 Introduction
Recently, Large Language Models (LLMs)(Touvron et al., 2023; Bubeck et al., 2023) demonstrate remarkable capabilities across various tasks, significantly improving the convenience of daily work and life. However, their large parameters and computational demands pose significant challenges for deployment. This makes quantization (Frantar et al., 2022; Lin et al., 2023; Shao et al., 2023) a crucial technology for reducing memory usage and speeding up inference (Yuan et al., 2024).
@@
Despite advancements, large outliers in LLMs activations can lead to significant quantization errors and accuracy loss. Many current methods address this by focusing on alleviating channel-wise outliers (Dettmers et al., 2022) through techniques like channel-wise scaling (Xiao et al., 2023a; Shao et al., 2023; Wei et al., 2023a), mixed-precision quantization (Dettmers et al., 2022; Zhao et al., 2023), Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024a), and channel-level assembly (Liu et al., 2023). However, activations of LLMs include not only channel-wise but also token-wise outliers. For example, Figure 1 (a) shows that some tokens, can be termed as outlier tokens, have extreme values exceeding 1,000, making it impractical to share quantization scales between outlier and normal tokens. The current leading method, Hadamard rotation (Ashkboos et al., 2024b), redistributes outlier values across all channels, reducing the maximum value in outlier tokens from over 1,000 to about 15 (see Figure 1 (b)). Nevertheless, the magnitude of outlier tokens remains hundreds of times greater than that of normal tokens, still suffering significant performance degradation when sharing quantization scales across different tokens.
Due to such dramatic discrepancies between normal and outlier tokens, previous quantization methods have to rely on per-token dynamic quantization to adjust quantization scales on-the-fly for each token. While per-token dynamic quantization adapts better to distribution changes, it faces more computational effort (Xiao et al., 2023a) and less compatible with operator fusion (Nagel et al., 2021) than per-tensor static quantization which use a fixed quantization parameter for all token. This leads to an important question: Can we eliminate token-wise outliers to enhance the precision of efficient per-tensor static quantization?
@@
Experiments demonstrate that, without any fine-tuning, PrefixQuant achieves comparable or better performance than previous per-token dynamic quantization methods (Ashkboos et al., 2024b; Xiao et al., 2023a; Lin et al., 2024b) using coarser per-tensor static quantization. Furthermore, fine-tuning significantly enhances PrefixQuant’s performance. For example, PrefixQuant with fine-tuning achieves a 7.43 WikiText2 (Merity et al., 2016) perplexity and 71.08% average accuracy across five common-sense reasoning tasks in W4A4KV4 Llama-3-8B, significantly outperforming previous QuaRot (Ashkboos et al., 2024b) with perplexity benefit and points accuracy. To the best of our knowledge, PrefixQuant is the first to outperform previous per-token dynamic quantization methods (Ashkboos et al., 2024b; Xiao et al., 2023a; Lin et al., 2024b) using coarse per-tensor static quantization. We also benchmark the end-to-end inference of W4A4 quantization, where PrefixQuant achieves a to speedup over FP16 models, and surpasses QuaRot models by to . We hope PrefixQuant inspires future developments in LLM compression.
2 Related Works
In this section, we discuss works related to outliers in LLMs, including quantization methods that enhance performance by eliminating activation outliers.
Channel-Wise Outliers. Dettmers et al. (2022) identifies that outliers in activation consistently occur in the same channels across different input tokens and proposes isolating these outlier channels with 16-bit precision. Other works, such as Atom (Zhao et al., 2023) and QUIK (Ashkboos et al., 2023), follow a similar mixed-precision approach to handle outliers. Instead of introducing mixed-precision matrix manipulation, which lacks native hardware support, another line of work addresses outliers through mathematically equivalent transformations. For example, SmoothQuant (Xiao et al., 2023a), OmniQuant (Shao et al., 2023), and Outlier Suppression (Wei et al., 2022; 2023b) mitigate outliers by scaling activations to weights on a channel-wise basis. QLLM (Liu et al., 2023) reduces outlier values by dividing each outlier channel into multiple sub-channels. Recently, QuaRot (Ashkboos et al., 2024b) proposed a simple and effective method, random Hadamard rotation, to redistribute outliers across all channels. Building on QuaRot, SpinQuant (Liu et al., 2024a) suggests training the orthogonal matrix instead of using a random Hadamard matrix to further enhance performance. DuQuant (Lin et al., 2024a) leverages channel permutation to evenly distribute outlier to each block and uses block-rotation to smoothen outliers. Although these methods significantly improve activation quantization performance, they all rely on fine-grained per-token dynamic quantization, which incurs additional overhead to manage token-wise fluctuations.
Token-Wise Outliers. The SoftMax function used in the self-attention mechanism naturally prevents producing zero attention scores. As a result, the model tends to assign unnecessary scores to special tokens, leading to token-wise outliers (or termed as massive activation) (Sun et al., 2024; Xiao et al., 2023b). Based on this, StreamingLLM (Xiao et al., 2023b) and LM-infinite (Han et al., 2023) support infinite sequences by retaining the initial token. Unlike StreamingLLM and LM-infinite, which simply preserve initial tokens in the KV-cache for long-context generation, our PrefixQuant carefully selects prefixed tokens in the KV-cache to isolate outliers for quantization. Some studies explore eliminating outliers with training techniques. For example, Bondarenko et al. (2024) allows SoftMax to produce zero values, and Sun et al. (2024) shows that adding attention bias in the KV cache during training can effectively reduce outliers. Our PrefixQuant efficiently isolates outlier tokens without needing retraining. The works closest to our approach are QFeP (Yang et al., 2024) and CushionCache (Son et al., 2024), which also set prefixed tokens in the KV cache. However, CushionCache (Son et al., 2024) takes 12 hours to find the prefixed tokens for Llama-3-8B through a greedy search, while our method completes this process in 12 seconds. QFeP (Yang et al., 2024) fixes the outlier token number for all models at 3, which lacks flexibility. Additionally, both QFeP and CushionCache suffer significant performance degradation when using per-tensor static quantization instead of per-token dynamic quantization. Our PrefixQuant is the first to make per-tensor static quantization outperform per-token dynamic quantization.
3 Preliminaries
Quantization in LLMs involves weight, activation, and KV cache quantization. Weight quantization (Chen et al., 2024a) and KV cache quantization (Liu et al., 2024b) reduce memory usage and speed up memory-bound computations (Yuan et al., 2024). Combining weight and activation quantization enables low-bit matrix manipulation to accelerate computation-bound tasks (Yuan et al., 2024). Specifically, the symmetric quantization process is:
where denotes rounding operation, is the target bit number, and are the quantized integer and full-precision activation, respectively. is the step size. Full precision weight can also be quantized into and similarly. Then, full-precision matrix manipulation transfer into efficient low-bit matrix manipulation:
Granularity. Finer granularity in quantization results in more overhead but less information loss. Per-tensor quantization shares across the entire tensor. Per-channel quantization of weight and per-token quantization of activation means is shared within each row of the tensor.
Dynamic and Static. Activation quantization divides into dynamic and static quantization based on how quantization parameters are calculated. Specifically, dynamic quantization calculates during inference, offering better adaptability to different distributions. In contrast, static quantization precomputes and in Eq.(2) offline, leading to more efficient inference and more feasible operator fusion (Nagel et al., 2021). Table 8(b) shows that the overhead of static quantization is nearly lower than dynamic quantization. Additionally, we initialize both and through grid search (Lin et al., 2023; Gong et al., 2024) on a small calibration dataset for all experiments with static quantization.
Hadamard Rotation. Random Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024a) addresses channel-wise outliers. Our method focus on removing token-wise outliers. Therefore, We build our method upon the Hadamard rotation technique, and the detailed is provided in Sec. C.
4 Difficulty of Static Quantization
Both channel-wise and token-wise outliers can cause information loss during quantization. While channel-wise outliers have been thoroughly explored and addressed in prior research (Ashkboos et al., 2024b), this discussion focuses on token-wise outliers, which occur within specific tokens.
Let represent the token sequence, with tokens and a dimension size of . We calculate token-wise maximum values , indicating the maximum value of each token. Per-tensor static quantization uses one pre-computed scale for all tokens. If the token-wise maximum values vary significantly across tokens, this can lead to substantial information loss after per-tensor static quantization. To analyze the distribution of token-wise maximum values and understand the challenges for per-tensor static quantization, we define top-1, median, and min-1 as the largest, median, and smallest values of , respectively. We then measure discrepancies using the ratios and . Specifically, a larger indicates upper outliers, while a larger represents lower outliers. Both ratios highlight the variability in . Specifically, we identify the following patterns that motivate our method.
1. Upper Outlier Tokens in Inputs. As shown in Figure 2(a), the input activation of down_proj layers exhibits significant outliers with . Although Hadamard rotation (Figure 2(b)) reduces the ratio to 478, it remains impractical to share a quantization scaling factor across tokens due to the large gap in maximum values.
2. Lower Outlier Tokens in //. We also investigate the distribution of // within the self-attention mechanism. We only quantize the KV cache for fair comparisons with previous works (Ashkboos et al., 2024b; Lin et al., 2024b). However, quantization of is also crucial, as used in FlashAttention-3 (Shah et al., 2024). In Figure 3, // display a different outlier pattern than the inputs of linear layers, with some tokens having extremely small magnitudes instead of large ones. Specifically, / have , but . Additionally, as shown in Figure 3(b), Hadamard rotation has no effect on these outliers.
[…]
Previous works (Ashkboos et al., 2024b; Lin et al., 2024b; Liu et al., 2024b) take per-token dynamic quantization for inputs of linear layers and KV cache to deal with outlier tokens. In this paper, we focus on eliminating outlier tokens to facilitate per-tensor static quantization.
5 PrefixQuant
As shown in Figure 5, we propose prefixing outlier tokens in the KV cache to improve the performance of more efficient per-tensor static quantization, instead of using costly per-token dynamic quantization. Section 5.1 explains how to find these prefixed outliers. Section 5.2 introduces block-wise fine-tuning to further enhance performance.
@@ 5.1 Prefixed Outliers
Definition of Outlier Token. Given that both upper outlier tokens in the inputs of the linear layer and lower outlier tokens in // are same tokens, we choose to identify outlier tokens using the upper outliers in the inputs of the down_proj layers due to the outlier in down_proj is more highlight and easier to be detected. Given token-wise maximum values , which represents the maximum values of each token. Then, outlier token in the -th index of token sequence is identified when the ratio of their maximum values to the median of all maximum values exceeds a threshold :
where is the maximum value of the -th token, denotes the function to find the median value from the vector, and the threshold is empirically set to in our experiments.
Number of Outlier Tokens. We determine the number of outlier tokens by calculating the average number of outlier tokens in a small calibration dataset. Specifically, we compute the average outlier token count for each transformer block according to Eq (3), where is the total number of transformer blocks. Since outlier tokens are nearly consistent across layers that contain them, we set the number of outlier tokens as .
Which Tokens to Prefix? Outlier tokens act as attention sinks (Xiao et al., 2023b), occupying only a few tokens () to help the attention mechanism do nothing (Bondarenko et al., 2024; Sun et al., 2024). Given the outlier token number , we find that prefixing the top- high-frequency***The frequencies are calculated without considering initial token. outlier tokens and the special ‘[BOS]‘ token can successfully constrains the outliers in prefixed tokens as shown in Figure 4(c). For special models (such as Llama-3-8B and Qwen-2-7B) with outlier tokens only in the initial tokens, we simply set the prefix token as “[BOS]”. The detailed prefixed tokens for different models are illustrate in Table 1. Considering the auto-regressive inference pipeline of LLMs, we store these prefix tokens in the KV cache to prevent generating new outlier tokens during inference. As shown in Figure 2(c) and Figure 3(c), prefixing outliers in the KV cache reduces the ratio of down_proj inputs from 476 to 2.7 and the ratio of / from to .
| PPL | FP16 | W16A4KV16 (static) | W16A16KV4 (static) | ||||
| original | + rotation | + prefixed | original | + rotation | + prefixed | ||
| Llama-2-7B | 5.47 | 3024.77 | 17.95 | 5.91 | 6.46 | 5.95 | 5.56 |
| Llama-3-8B | 6.14 | 1816.57 | 22.14 | 7.23 | 7.37 | 8.12 | 6.30 |
Quantitative Analysis. Table 2 presents separate performance of static quantization on input activation and KV cache quantization. We can find that the model suffers significant performance degradation with static quantization becuase of outlier tokens. For example, in Llama-3-8B, WikiText2 perplexity increases from 6.14 to 22.14 with 4-bit per-tensor activation quantization and from 6.14 to 8.12 with 4-bit per-head static KV cache quantization even with Hadamard rotation. After further setting prefix outliers in the KV cache, performance significantly improves: perplexity of 4-bit per-tensor activation decreases to 7.23 and perplexity of 4-bit per-head static KV cache quantization decreases to 6.30, demonstrating the effectiveness of prefixed outlier tokens for static quantization.

### v2

# PrefixQuant: Eliminating Outliers by Prefixed Tokens for Large Language Models Quantization

@@ Abstract
Existing weight-activation quantization methods for Large Language Models (LLMs) primarily address channel-wise outliers but often neglect token-wise outliers, which limits the accuracy of quantized models. In this work, we propose PrefixQuant, a novel quantization method that achieves state-of-the-art performance across various precision levels (W4A4KV4 and W4A8KV4) and granularities (dynamic and static quantization) by effectively isolating token-wise outliers. First, PrefixQuant eliminates token-wise outliers by prefixing outlier tokens in the KV cache, a process that is training-free and highly efficient (e.g., 1 minutes for Llama-3-70B). Second, PrefixQuant introduces new trainable parameters for block-wise training to compensate for quantization error. Our experiments show that PrefixQuant significantly outperforms existing dynamic quantization methods, even under coarser static quantization settings. For instance, PrefixQuant achieves an average accuracy improvement of and points over SpinQuant (dynamic quantization) on five zero-shot reasoning tasks under dynamic and static quantization settings, respectively, on W4A4KV4 Llama-3-8B. Additionally, we demonstrate up to prefilling speedup and decoding speedup for LLMs using W4A4 PrefixQuant. Our code is available at https://github.com/ChenMnZ/PrefixQuant.
@@ 1 Introduction
Recently, Large Language Models (LLMs)(Touvron et al., 2023; Bubeck et al., 2023) demonstrate remarkable capabilities across various tasks. However, their large parameters and computational demands pose significant challenges for deployment. This makes quantization (Frantar et al., 2022; Lin et al., 2023; Shao et al., 2023) a crucial technology for reducing memory usage and speeding up inference (Yuan et al., 2024).
@@
Despite advancements, outliers in LLMs activations can lead to significant quantization errors and accuracy degeneration. Many current methods address this by focusing on alleviating channel-wise outliers (Dettmers et al., 2022) through techniques like channel-wise scaling (Xiao et al., 2023a; Shao et al., 2023; Wei et al., 2023a), mixed-precision quantization (Dettmers et al., 2022; Zhao et al., 2023), Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024b), or channel-level assembly (Liu et al., 2023). However, activations of LLMs include not only channel-wise outlier but also some massive activation (Sun et al., 2024) only occur in a few tokens, and can be termed as token-wise outliers. For example, Figure 1 shows that 2 outlier tokens among 2048 tokens contribute 94.7% of the quantization error. Figure 2(a) provides a more detailed analysis, revealing that a few tokens have extreme values exceeding 1,000, resulting in a quantization error of . The current state-of-the-art method, Hadamard rotation(Ashkboos et al., 2024b), redistributes outlier values across all channels, reducing the maximum value of outlier tokens from over 1,000 to approximately 15 (see Figure 2 (b)). However, the magnitude of outlier tokens remains hundreds of times larger than normal tokens, leading to a quantization error of .
In this paper, we propose PrefixQuant, an efficient method to isolate token-wise outliers for more accurate quantization. PrefixQuant is based on a key observation: Prefixing high-frequency outlier tokens at the beginning of the input sequence constrains token-wise outliers to only occur in the prefixed tokens. Since prefixed tokens remain consistent across all inputs, PrefixQuant performs offline prefilling of these tokens and stores their KV cache. This stored KV cache can then be reused for all inputs, effectively avoiding token-wise outliers during the forward pass. Furthermore, the detection of prefixed tokens is efficient and does not require any retraining, unlike prior methods (Sun et al., 2024; Bondarenko et al., 2024). For example, this process completes in just 12 seconds for Llama-2-7B. As shown in Figure 2(c), PrefixQuant effectively eliminates outliers and reduces the quantization error from 7.88 (using QuaRot) to 0.04. Additionally, we introduce a block-wise fine-tuning (Shao et al., 2023; Chen et al., 2024a) to compensate for quantization error by jointly training weights and quantization parameters. For static activation quantization, the quantization parameters are inherently trainable. However, dynamic activation quantization lacks trainable quantization parameters. To address this, we propose learnable activation clipping to enable training for dynamic activation quantization.
Since PrefixQuant is compatible with various quantization schemes, we introduce two settings for PrefixQuant (see Table 1): O1 for dynamic quantization and O2 for static quantization. Experiments show that PrefixQuant significantly outperforms existing methods (Ashkboos et al., 2024b; Xiao et al., 2023a; Lin et al., 2024b) under the same dynamic quantization setting. Furthermore, PrefixQuant even surpasses prior dynamic quantization methods while utilizing the more efficient static quantization. For instance, PrefixQuant achieves an average accuracy improvement of and points over SpinQuant (dynamic quantization) on five zero-shot reasoning tasks under dynamic and static quantization settings, respectively, on W4A4KV4 Llama-3-8B. To the best of our knowledge, PrefixQuant is the first method to surpass prior per-token dynamic quantization methods (Ashkboos et al., 2024b; Xiao et al., 2023a; Lin et al., 2024b) using the coarser per-tensor static quantization. We also benchmark the end-to-end inference of W4A4 quantization, where PrefixQuant achieves a prefilling speedup and a decoding speedup compared to FP16 models. We hope PrefixQuant inspires future developments in LLM compression.
| Method | Weight | Activation | KV Cache |
| SmoothQuant | per-channel | per-token dynamic | per-token dynamic |
| Atom | group-wise | group-wise dynamic | group-wise dynamic |
| QoQ;QuaRot;SpinQuant;SpinQuant | per-channel | per-token dynamic | group-wise dynamic |
| PrefixQuant-O1 | per-channel | per-token dynamic | group-wise dynamic |
| PrefixQuant-O2 | per-channel | per-tensor static | per-head static |
2 Related Works
This section discusses works related to enhancing quantization accuracy by eliminating activation outliers.
Channel-Wise Outliers. Activation outliers often recur in the same channels across tokens. (Dettmers et al., 2022) addresses this by isolating outlier channels with 16-bit precision, while Atom (Zhao et al., 2023) and QUIK (Ashkboos et al., 2023) adopt similar mixed-precision strategies. Other methods, like SmoothQuant (Xiao et al., 2023a), OmniQuant (Shao et al., 2023), and Outlier Suppression (Wei et al., 2022, 2023b), scale activations to weights on a channel-wise basis. QLLM (Liu et al., 2023) splits outlier channels into sub-channels, and QuaRot (Ashkboos et al., 2024b) redistributes outliers using random Hadamard rotation, later improved by SpinQuant (Liu et al., 2024b), which trains the orthogonal matrix. In contrast, our work focuses on token-wise outliers and is orthogonal to these channel-wise methods.
Token-Wise Outliers. The SoftMax function in self-attention prevents zero attention scores, causing unnecessary scores for special tokens and leading to token-wise outliers (Sun et al., 2024; Xiao et al., 2023b; Gu et al., 2024). StreamingLLM (Xiao et al., 2023b) and LM-infinite (Han et al., 2023) retain initial tokens for long-context generation, while our PrefixQuant isolates outliers by carefully selecting prefixed tokens in the KV-cache for quantization. Unlike training-based methods (Bondarenko et al., 2024; Sun et al., 2024) that modify SoftMax behavior or add attention bias, our PrefixQuant isolates outliers without retraining. Closest works, like QFeP (Yang et al., 2024) and CushionCache (Son et al., 2024), rely on costly grid searches (e.g., 12 hours for Llama-3-8B), while PrefixQuant completes this in 12 seconds. Furthermore, unlike prior works (Sun et al., 2024; Son et al., 2024) focusing on large-value outliers, PrefixQuant also identifies extremely small-value outliers in self-attention queries and keys.
3 Preliminaries
Quantization in LLMs involves weight, activation, and KV cache quantization. Weight quantization (Chen et al., 2024a) and KV cache quantization (Liu et al., 2024a) reduce memory usage and speed up memory-bound computations (Yuan et al., 2024). Combining weight and activation quantization enables low-bit matrix manipulation to accelerate computation-bound tasks (Yuan et al., 2024). Specifically, the quantization process is:
where denotes rounding operation, is the target bit number, and are the quantized integer and full-precision tensor, respectively. and are quantization parameters, for the step size and zero values, respectively. and are clipping factors.
Dynamic and Static. Activation quantization is divided into dynamic and static quantization based on how quantization parameters are calculated. Specifically, dynamic quantization calculates and online during inference, offering better adaptability to different distributions. In contrast, static quantization precomputes and offline through calibration datasets, leading to more efficient inference and more feasible operator fusion (Nagel et al., 2021).
Initialization of Quantization Parameters. The classical approach uses max–min initialization, where both and are set to . To better balance clipping error and rounding error (Lin et al., 2023, 2024b), we initialize and using MSE-based grid search for both weight and activation quantization in our experiments. Specifically, for per-token dynamic quantization, and are shared across all tokens within the same layer. For per-tensor static quantization, we directly perform a grid search for the quantization parameters and instead of optimizing the clipping factors.
Hadamard Rotation. Random Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024b) addresses channel-wise outliers. Our method focus on removing token-wise outliers. Therefore, We build our method upon the Hadamard rotation technique, and the detailed is provided in Sec. B.
4 PrefixQuant
In this section, we present the proposed PrefixQuant methods. Sec.4.1 describes the characteristics of outlier tokens in LLMs. Sec.4.2 proposes a solution to isolate these outlier tokens, creating a more quantization-friendly distribution. Finally, Sec.4.3 introduces block-wise optimization to further reduce quantization error.
4.1 Deep Exploration of Outlier Tokens
Both channel-wise and token-wise outliers can cause significant quantization error. While channel-wise outliers have been thoroughly explored and addressed in prior research (Ashkboos et al., 2024b), this discussion focuses on token-wise outliers.
Definition of Outlier Token. Let represent the absolute values of a token sequence, where is the number of tokens and is the dimension size. We compute the token-wise maximum values , where each element indicates the maximum value of the -th token. The outlier degree of a token is then measured by comparing its maximum value to the median of :
Then, the -th token is classified as an outlier if deviates significantly from 1 (i.e., either much larger or much smaller than 1). Specifically, we define an upper outlier token when and a lower outlier token when . In our experiments, we set and .
Visualization of Outlier Tokens. To better illustrate the outlier degree, we further define and as the maximum and maximum across different layers, respectively. Following the definition of in Eq. (3), a larger indicates the presence of extreme upper outliers, while a larger reflects the presence of extreme lower outliers. Specifically, we identify the following outlier tokens:
1) Upper outlier tokens in inputs of down_proj layers and outputs of transformer blocks. As shown in Figure 3(I.a), the input activations of the down_proj layers exhibit significant upper outliers, with . Although Hadamard rotation (Figure 3(I.b)) reduces this ratio to 461, it still indicates a large gap compared to normal tokens. A similar phenomenon is observed in the outputs of transformer blocks, as shown in Figure 3(II). These outliers not only lead to larger quantization errors but also cause instability during block-wise fine-tuning.
2) Lower outlier tokens in /. In Figure 3(II.a), displays a distinct outlier pattern compared to the inputs of linear layers. Instead of having large magnitudes, some tokens exhibit extremely small values. Specifically, has , but . Furthermore, as shown in Figure 3(II.b), Hadamard rotation does not mitigate these lower outliers. Similar lower outliers are also observed in , as shown in Figure 10.
Additionally, we observe that both upper outlier tokens and lower outlier tokens correspond to tokens at the same position in the sequence, but they exhibit different patterns in different modules. Therefore, we focus on analyzing upper outlier tokens due to their stronger prominence and ease of detection.
Characters of outlier tokens. We further investigate the characteristics of these outlier tokens, including the number of outlier tokens in an input sequence, their positions, and their content (text):
Number: We determine the number of outlier tokens in a small calibration dataset. Specifically, we compute the average outlier token count for each transformer block according to compare Eq (3) with the outlier threshold , where is the total number of transformer blocks. Since outlier tokens are nearly consistent across layers that contain them, we simply set the number of outlier tokens as . Consistent with Massive Attention (Sun et al., 2024), we find that outlier tokens appear in only a small fraction of positions (e.g. 2 for Llama-2-7B) within the input sequence, as shown in Figure 4(a).
Position: We observe that the initial tokens are outlier tokens across almost all models, aligning with findings on attention sinks (Xiao et al., 2023b). Additionally, Figure 4(c) shows that, apart from the initial tokens, some other tokens near the beginning of the sequence are also outlier tokens. Unlike outlier channels, which occur at fixed channel indices (Dettmers et al., 2022), the positions of outlier tokens depend on the input sequence and vary significantly. As a result, it is not feasible to identify outlier tokens offline for mixed-precision quantization.
@@ 4.2 Prefixed Outliers
Given that the number of outlier tokens is limited and they typically occur at the beginning of the input sequence, we propose a method to prefix high-frequency outlier tokens in the input sequence. This approach constrains outlier tokens to the prefixed tokens. Furthermore, these prefixed tokens can be directly stored in the KV cache, enabling more efficient computation, as shown in Figure 2.
@@
What token is added as a prefix. To determine which tokens to add as a prefix, we firstly analyze the number of outlier tokens . We find that prefixing the top- high-frequency111The frequencies are calculated excluding the initial token. outlier tokens successfully constrains outliers to the prefixed tokens, as illustrated in Figure 4(d). For special cases, such as models like Llama-3-8B and Qwen-2-7B where outlier tokens only appear as initial tokens, we set the prefix token to ”[BOS]”. Additionally, for consistency, we also include ”[BOS]” as the last prefixed token for all models. The detailed prefixed tokens used for different models are listed in Table 5.
Computation of prefixed tokens in KV cache. In the auto-regressive inference pipeline of LLMs, we directly store these prefixed tokens in the KV cache to prevent new outlier tokens from being generated during inference. Specifically, given the input query, key, and value matrices , the self-attention mechanism with prefixed tokens in the KV cache is formulated as:
Here, are the prefixed tokens stored in the KV cache. We compute and during a one-time prefilling process using the full-precision model. These prefixed tokens are then stored in the KV cache and reused during inference by quantized models. Notably, the prefixed tokens in the KV cache remain in full precision, even when used with quantized models.
Distribution changing after setting prefixed tokens. As shown in Figure 3, prefixing outlier tokens in the KV cache significantly improves the distribution. Specifically, the ratio of the down_proj inputs decreases from 461 to 2.4 and the ratio of / decreases from to .